A net force of 500 N gives an object an acceleration of 4.00 meter per s^2 a.) What net force is needed to give the object an acceleration of 12 meter per s^2? b.) What net force is required to keep the object moving at a constant velocity?
1
Expert's answer
2019-12-09T09:05:08-0500

a)


"m=\frac{F_1}{a_1}=\frac{500}{4}=125\ kg"

"F_2=ma_2=125(12)=1500\ N"

b)


"F=0 \ N"